OpenStudy (chey):

x-c/2=d soolve for x *show work*

OpenStudy (anonymous):

To solve for x, you need gather all other terms on the other side of the equation.

OpenStudy (lalaly):

\[\frac{x-c}{2}=d\]\[\cancel{2} \times \frac{x-c}{\cancel{2}}= d \times 2\]\[x-c=2d\]\[x=2d+c\]